OpenStudy (anonymous):

Find all the complex roots. Leave in polar form with the arguments in degrees. The complex cube roots of 1+i? Simplify the answer, including radicals. Type an exact answer, including radicals as needed.

OpenStudy (anonymous):

\[z_0=?(\cos?^o+isin?^o)\]

OpenStudy (anonymous):

Use DeMoivre's Theorem

OpenStudy (anonymous):

\[(\cos \theta + i \sin \theta)^{n} = \cos n \theta + \sin n \theta\]

OpenStudy (anonymous):

I know the theorem.. but, I have no idea how to apply it.

OpenStudy (phi):